# Unbiased Learning to Rank Meets Reality: Lessons from Baidu’s Large-Scale Search Dataset

### v1

@@ 8. Conclusion
Our main conclusion, however, is that while we have observed substantial improvements relating to the choice of query-document representations and ranking loss (e.g., pointwise or listwise), conventional unbiased learning-to-rank techniques do not bring any clear benefits when applied to the Baidu-ULTR dataset. Our findings confirm the work by Zou et al. (2022), even though we used different dataset features and preprocessing, model implementations, and performed extensive hyperparameter tuning. Further, we observed a divergence between click-based and annotations-based objectives, as all click-based approaches were outperformed on annotation-based metrics by simple baselines such as BM25, even when BM25 was a model input. We believe these results call for more research into the conditions for success and failure of unbiased learning-to-rank and click-based approaches as a whole.
@@ Acknowledgements.
We thank Lixin Zou for his valuable insights and clarifications on the original Baidu work, as well as Jiaxin Mao and Zechun Niu for sharing their learnings from NTCIR 17. This research was supported by the Mercury Machine Learning Lab, created by TU Delft, the University of Amsterdam, and Booking.com. Maarten de Rijke was supported by the Hybrid Intelligence Center, a 10-year program funded by the Dutch Ministry of Education, Culture and Science through the Netherlands Organisation for Scientific Research, https://hybrid-intelligence-centre.nl. All content represents the opinion of the authors, which is not necessarily shared or endorsed by their respective employers and/or sponsors.

### v2

@@ 8. Conclusion
Our main conclusion, however, is that while we have observed substantial improvements relating to the choice of query-document representations and ranking loss (e.g., pointwise or listwise), conventional unbiased learning-to-rank techniques do not bring clear improvements in ranking performance on the Baidu-ULTR dataset. Our findings confirm the work by Zou et al. (2022), even though we used different dataset features and preprocessing, model implementations, and performed extensive hyperparameter tuning. Further, we observed a divergence between click-based and annotations-based objectives, as all click-based approaches were outperformed on annotation-based metrics by simple baselines such as BM25, even when BM25 was a model input. We believe these results call for more research into the conditions for success and failure of unbiased learning-to-rank and click-based approaches as a whole.
@@ Acknowledgements.
We are grateful to Lixin Zou for his valuable insights and clarifications on the original Baidu work, as well as to Jiaxin Mao and Zechun Niu for sharing their learnings from NTCIR 17. This research was supported by the Mercury Machine Learning Lab, created by TU Delft, the University of Amsterdam, and funded by Booking.com. Maarten de Rijke was supported by the Hybrid Intelligence Center, a 10-year program funded by the Dutch Ministry of Education, Culture and Science through the Dutch Research Council, https://hybrid-intelligence-centre.nl. All content represents the opinion of the authors, which is not necessarily shared or endorsed by their respective employers and/or sponsors.